# Renovating Names in Open-Vocabulary Segmentation Benchmarks

### v1

@@ 4.3 Training with renovated names
Setup. In this experiment, we investigate the benefit of using renovated names for training open-vocabulary models. Specifically, we train an FC-CLIP model on MS COCO with our renovated segment-matched names for the training set. We also train two other FC-CLIP models using the baseline names.
For a fair comparison of these models trained with different sets of names, during inference, we merge the three name sets to form a single class name set as the text prompt. For RENOVATE, the names for each class are simply the set of unique names for all the instances in that class. For evaluation, we group the name predictions based on the original class division. In this way, even when using more fine-grained prompts, we still compute the evaluation metrics for each original class. It’s important to note that our evaluation methodology aligns with those employed in previous studies [11, 48, 52] since the original and OpenSeg names also similarly accommodate multiple naming options for their classes. Since we compute metrics on the same class divisions as the original datasets, we only report standard metrics (PQ, AP, and mIoU) for comparing different models. We observe similar evaluation results when using open metrics. We detail them in the supplements.
@@
RENOVATE 56.62 45.50 65.48 27.98 17.94 37.50 46.10 27.99 61.25
@@
Results. Our results in Tab. 3 demonstrate that using RENOVATE names for training improves segmentation on both source dataset and target datasets. For example, the model trained with RENOVATE names obtains over 2% PQ on MS COCO, near 2% mIoU on ADE20K and around 3% AP and mIoU on Cityscapes. In addition, we notice that the mIoU metric typically gets improved more with our names than AP. This is consistent with our empirical findings that “stuff names” are typically more vague and thus have improvement potential than “thing names”. These results suggest that more precise and descriptive training names enhance the model generalization to both known and novel categories, potentially by better retaining the pre-trained vision-language alignment of the CLIP backbone during fine-tuning.
4.4 Automating name quality evaluation
Setup. We further leverage pre-trained open-vocabulary models as evaluators of names, automating the name quality evaluation process. Instead of involving humans, we prompt open-vocabulary segmentation models with different names. Names are then considered to be better if they help the model achieve better open-vocabulary segmentation. During inference, the text prompt consists of all the unique names from the corresponding name set. For evaluation, we adopt the same evaluation protocol as in Sec. 4.3 to compute metrics (PQ, AP, mIoU) on the same class divisions as the original datasets.
Panoptic Semantic
ADE-150 [57] Cityscapes [7] ADE-847 [57] PC-459 [27]
Model Test names PQ AP mIoU PQ AP mIoU mIoU mIoU
ODISE [48] Original names 21.88 13.94 29.16 39.72 27.73 49.60 10.75 16.97
OpenSeg names 21.63 13.95 28.88 43.26 28.45 54.53 11.15 17.81
RENOVATE 23.69 14.38 31.64 43.61 28.38 57.42 11.82 18.35
MasQCLIP [51] Original names 23.46 12.80 30.32 33.78 18.11 45.35 10.82 18.21
OpenSeg names 23.70 12.84 31.17 35.05 17.79 46.46 10.95 18.36
RENOVATE 25.00 12.93 32.30 35.63 18.19 50.07 11.39 18.85
FC-CLIP [52] Original names 24.30 16.79 32.14 39.42 22.76 53.08 13.33 18.32
OpenSeg names 25.35 17.30 33.06 43.68 26.93 56.15 13.78 19.22
RENOVATE 27.90 17.88 37.05 45.90 29.79 62.55 15.63 20.74
Table 4: Open-vocabulary segmentation evaluation. We use ODISE, MasQCLIP, and FC-CLIP pre-trained on COCO [22] and evaluate both panoptic and semantic segmentation performance. Our results show that RENOVATE class names are of high quality since they help boost the performance of open-vocabulary models.
Results. In Tab. 4, we experiment with three open-vocabulary segmentation models, i.e., MasQCLIP [51], ODISE [48], and FC-CLIP [52], all pre-trained on the MS COCO dataset. RENOVATE names consistently outperform the original and OpenSeg names across ADE20K, Cityscapes, and PASCAL Context datasets. Notably, FC-CLIP with RENOVATE names attains 15.6% and 16.4% relative PQ improvement from the original names on ADE20K and Cityscapes, respectively, setting new state-of-the-art performances on both benchmarks. This suggests that RENOVATE names align more accurately with the visual segments, enhancing segmentation performance. Note that some numbers in Tab. 4 are even better than when we train models with renovated names (Tab. 3). That is because the test names in Tab. 3 are of larger size with mixed name qualities from all three name sets, presenting a greater challenge. Despite this, when tested using the same names, as in Tab. 3, models trained with RENOVATE names are stronger than the others.
In addition, we observe again that mIoU is typically improved more with RENOVATE names than AP, further confirming our hypothesis that “stuff” class names have more improvement room. Moreover, we note that RENOVATE improves the performance of ODISE, which employs a Stable Diffusion [35] backbone instead of CLIP. This indicates the general applicability of RENOVATE names across different VLM backbones.
4.5 Ablation Analysis on the Renaming Pipeline
We further perform an ablation study using the same automated name quality evaluation as in Sec. 4.4. In Tab. 6, we first analyze the impact of context names for candidate generation. Our results indicate significant performance improvements when context information is provided, as opposed to the “None” approach, which relies only on original names. This highlights the effectiveness of leveraging contextual insights for generating superior candidate names. In comparison, CaSED outperforms both BLIP2 and RAM, achieving a greater boost in performance through its retrieval-based approach and the utilization of expansive external image-caption datasets, underlining the benefits of contextual enrichment in name generation.
Table 5: Ablation on the context name sources.
ADE20K Context name source PQ AP mIoU None 25.50 16.53 33.80 Caption nouns (BLIP2 [19]) 25.89 17.36 34.74 Image tags (RAM [55]) 26.95 17.55 35.54 Caption nouns (CaSED [6]) 27.90 17.88 37.05
ADE20K Attn bias Query Classifier PQ AP mIoU gt mask learnable text 21.09 14.91 26.96 gt mask text learnable 25.65 16.33 32.67 rand gt mask text learnable 26.21 17.50 35.62 rand gt mask text w/ neg learnable 27.90 17.88 37.05
Table 5: Ablation on the context name sources.
Table 6: Ablation on renaming model architecture.
In Tab. 6, we analyze the design choices of our renaming model. We start by establishing that our approach using text queries is superior to a standard text classifier, which only interacts with the transformer decoder’s output query embeddings and treats the mask prediction branch in a class-agnostic way. In comparison, our text queries interact intimately with spatial details from image features and exploit the mask prediction branch, leading to more precise name-segment matching. We also observe performance benefits from randomly substituting ground-truth masks with predicted ones in attention biases (“rand gt mask”), which suggests training with increased reliance on text queries refines the matching quality. Furthermore, augmenting our model with negative name appending (“text w/ neg”) enhances its discriminative capacity, further boosting performance and solidifying the renaming model’s overall effectiveness.
5 Discussion and Outlook
In this work, we highlight the problems of naming in existing open-vocabulary segmentation benchmarks and introduce RENOVATE, a straightforward yet versatile framework to renovate dataset names. Our efforts underscore the benefits of renaming, evidenced by the improvements in upgraded benchmarks and models trained with renovated names. However, while RENOVATE helps uncover biases in models and datasets, it could inadvertently propagate biases from the foundational models into the new names. To mitigate potential negative societal impacts, we advocate for verification of renovated names in critical applications, as we have exemplified with our upgraded benchmarks.
As the first attempt to propose a generalized, principle-based renaming framework, we acknowledge that our exploration of design choices remains incomplete. For instance, alternative language models like Llama [44] or Gemini [43] have yet to be investigated for candidate name generation, and our architectural experiments have so far been confined to using CLIP as the backbone, leaving other VLMs like SigCLIP [53] unexplored. In the future, we aim to further refine our renaming method, exploring a wider array of options, and ultimately scale it to automate the enhancement of publicly available large-scale datasets.
@@ 5 Discussion and Outlook
  • In Appendix 0.A, we provide a brief overview of works that address name problems in open-vocabulary classification.
@@
  • In Appendix 0.B, we elaborate on the implementation details on the GPT-4 prompts for name generation, human verification process for benchmark upgrading, open metrics, human studies, and the renaming and pre-trained models.
  • In Appendix 0.C, we provide more experiment results on the upgraded benchmarks and on models trained with renovated names.
@@
  • In Appendix 0.D, we present more examples of context, candidate, and renovated names. We also show more examples of name selection with the IoU scores.
@@
  • In Appendix 0.E, we provide more qualitative examples demonstrating that RENOVATE can refine original names, correct wrong annotations, and uncover segments with shared semantic concepts across datasets.
Appendix 0.A A brief overview of name problems in open-vocabulary classification
@@
While OpenSeg [11] is the only prior work for the name problems in open-vocabulary segmentation, there are more research works on improving name quality for open-vocabulary classification. For example, [10] improve class names by including their parents and children from the WordNet hierarchy [26].  [6] propose to replace the fixed vocabulary list with nouns extracted from best-matching captions.  [24] improve the descriptiveness of the vocabulary by adding class descriptions. However, since the VLMs used by these models are all trained on image-text datasets, it is non-trivial to extend them to dense prediction tasks like segmentation. For the renaming strategy, our context name design and LLM-based generation approach ensure the diversity of our names with hierarchical concepts as well as descriptive contexts, without the use of an external word hierarchy as [26] or class descriptions as [24]. Our renaming model further improves the name quality based on the visual segments and is more specifically designed for the segmentation tasks.
Appendix 0.B More implementation details
0.B.1 GPT-4 prompts for name generation
@@
For the system message, we start the message with “You are a helpful assistant” following the examples in OpenAI Chat Completions API [1]. We then describe our problem setup, input format, and instructions with examples on name generation. For the user message, {original class name} and {context names} require inputs from each category. For our ablation experiments without context names (“None” in Tab. 6 in the main paper), we remove the descriptions about context names in the system message and do not require context names in the user message.
0.B.2 More details on the human verification process
We upgrade the benchmarks on ADE20K and Cityscapes with our renovated names. To ensure the high quality of names, we verify the names on all the segments in their validation set. As shown in Fig. 0(a), for each segment, we provide the model-selected name along with the other top 3 names ranked by the model, we also provide the rest of the candidate names in the “others” option. Therefore, the human verifier can quickly continue to the next segment if the selected name itself or another name in top 3 names matches the most with the segment. If none of the top 3 names fits well, humans need to choose from the other candidate names in “others”. In the latter case, it would be similar to the scenario where RENOVATE suggestions are not provided and humans have to choose from the whole un-ranked candidate name list. In Fig. 0(b), we compare the name verification speed with and without RENOVATE suggestions on 200 segments. As we can see, RENOVATE suggestions significantly speed up our verification process. We note that the human verifiers are provided only with candidate names from the same original class, which is typically the original ground-truth class. In the rare cases where our renaming model prefers names from classes other than the original groundtruth class, we first verify the class choices, then provide the human verifiers with the name suggestions from the verified class.
We further note that for all the segments verified, 68% of them are matched to the default, i.e., top 1, model-selected name, and 92% of them are matched to the top 3 RENOVATE suggestions, reflecting the high quality of RENOVATE names. In Fig. 0.B.2, we show typical cases when humans and models disagree, including segments that are too small to recognize and segments whose names cannot be inferred correctly due to insufficient visual cues. We note that such segments only consist a minority of all the segments, with inherent difficulty to name precisely.

### v2

@@ 5.2 Training with renovated names
Setup. We train FC-CLIP [25] models on MS COCO with our renovated segment-matched names and compare them with models trained using other name sets on COCO, ADE20K, and Cityscapes. Specifically, “OpenSeg” names [5] were manually curated by previous researchers. “Synonym names” are generated by prompting GPT-4 based on the original class names. “Candidate names” are the outputs of the initial step in our renaming process, using image captioning to provide context for prompting GPT-4. A key difference between these names and RENOVATE names is that the former uses a set of names for a class of segments, whereas RENOVATE names provide a name for each individual segment. For model evaluation, we follow prior practices [5, 24, 25] to group predictions based on the original class divisions. For a fair comparison, our test name set merges names from Original, OpenSeg, and RENOVATE names.
Results. Our results in Table 3 demonstrate that using RENOVATE names for training improves segmentation on both source dataset and target datasets. For example, compared to the Original names, the model trained with RENOVATE names improves by near 4 PQ on MS COCO and over 5% mIoU on ADE20K, showing significantly larger gains offered by OpenSeg names. This underscores the value of name refinement on per-segment level and indicates the high quality of RENOVATE names. We also note that models trained with Synonym names and Candidate names show inferior performance even to the Original names, reflecting the importance of high name quality.
@@
Synonym names 23.38 35.87 47.20 15.56 14.78 29.74 22.94 19.58 45.91
Candidate names 41.32 40.96 56.68 19.80 15.85 33.43 39.20 20.79 51.80
RENOVATE 56.62 45.50 65.48 27.98 17.94 38.50 46.10 27.99 61.25
@@
(a) MS COCO $\rightarrow$ ADE20K.
(b) MS COCO $\rightarrow$ MS COCO.
Figure 5: Data efficiency comparison between RENOVATE and original names.
Figure 6: Effectiveness of negative sampling.
The richer text information offered by RENOVATE names also improves training efficiency. In Fig. 6, we show that models trained with RENOVATE names can achieve comparable performance with significantly less data compared to those trained with the Original names. For instance, models trained with $10^{4}$ images using RENOVATE names can match or exceed the performance of models trained with $5\times 10^{4}$ images using the Original names. This corroborates previous findings that improving data quality can reduce the required quantity of training data [43, 44].
Ablation on the negative sampling. In Fig. 6, we study the effectiveness of negative sampling. Interestingly, we found that negative sampling is especially helpful to RENOVATE names while minorly impacting models trained with original names. This is because RENOVATE names have much more class names and the names are also more semantically close. Negative sampling effectively controls the number of classes used in cross entropy calculation and keeps the name embeddings relatively distinct, thus significantly improving training effectiveness for RENOVATE names.
5.3 Improving evaluation with renovated names
Setup. We evaluate pre-trained FC-CLIP [25], MasQCLIP [33], and ODISE [24] models on ADE20K and Cityscapes using three different name sets with both standard and open metrics. For open metric computation, we need the similarity between the ground-truth and predicted class names. For Original and OpenSeg names, which use a set of names for a class of segments, we use the averaged pairwise name similarity between classes. For RENOVATE names, we can directly use the pairwise name similarity between renovated names. Fig. 7 shows PQ evaluation of pre-trained models on ADE20K. Full evaluation results, including mAP and mIoU and results on Cityscapes are provided in Section B.3.
(a) FC-CLIP [25]
(b) MasQCLIP [33]
(c) ODISE [24]
Figure 7: Open-vocabulary evaluation on ADE20K with different names, metrics, and models.
Results. In Fig. 7, we first note that the discrepancy between standard and open metrics is significantly greater when evaluated using RENOVATE names. This indicates that many misclassification cases are indeed “benign”, i.e., when evaluated using more fine-grained names, the misclassified names turned out to be semantically close to the ground-truth names. For example, when an “area rug” is predicted to be a “floor rug” (RENOVATE names), it will be penalized much less than from “rug” to “floor” (original names); but if an “area rug” is predicted as “concrete floor”, the classification error will be even higher. This shows that current pre-trained models have learned relatively good semantic concepts, while still being sensitive to the text prompts at inference time.
(a) Per-category IoU on the ADE20K with 578 RENOVATE classes.
(b) Misclassification analysis.
Figure 8: RENOVATE names enable more fine-grained analysis on models. (a) Per-category IoU with highlighted top/bottom-5 RENOVATE names and selected names from the “person" class in ADE20K. (b) Confusion matrix of frequently misclassified RENOVATE names from “building" to “wall", showing misclassification proportions (numbers) and pairwise name similarity (color).
RENOVATE names also enable a more fine-grained analysis of the models. In Fig. 8(a), we present a per-category IoU analysis on RENOVATE names within the ADE20K dataset. We can see that large, well-defined objects such as “sky” and “girl” have high IoU scores, while small, deformable, and rare classes like “bollard” and “ceiling light” are more challenging for current models. Note that most of these names do not exist in the original benchmarks. In addition, our analysis reveals model biases, evidenced by the disparity in IoU scores between “man” and “woman” as well as “girl” and “boy”, suggesting an imbalance in the training dataset. In Fig. 8(b), we further illustrate how RENOVATE names facilitate a more detailed investigation of misclassifications. For example, among “building” segments misclassified as “wall”, 32.7% “churches” from the “building” class are predicted to be “brick wall” from the “wall” class. Additionally, we observe a strong correlation between name similarity and misclassifications. This showcases that our RENOVATE names can help identify problematic sub-categories within the original class that models still struggle with.
6 Conclusion and Limitations
In this work, we address the naming issues and show that renaming improves both model training and evaluation of open-vocabulary segmentation. While RENOVATE uncovers model biases, it could inadvertently propagate biases from the foundational models into the new names. To mitigate potential negative societal impacts, we advocate for verification of names in critical applications, as exemplified by our verification of names in evaluation sets. As the first attempt to propose a generalized renaming framework, we acknowledge that our exploration remains incomplete. In the future, we aim to further refine our method, exploring more design choices such as other model backbones [45], and scale it up to the publicly available large-scale datasets [46, 47].
@@ 6 Conclusion and Limitations
  • In Appendix A, More Literature Review, we provide a brief overview of works that address name problems in open-vocabulary classification.
  • In Appendix B, More Experiment Results, we provide more experiment results, including a human preference study on the name quality, additional ablation analysis on the renaming pipeline, full results on improving evaluation using renovated names and some more examples of using RENOVATE names to conduct fine-grained misclassification analysis.
@@
  • In Appendix C, More Implementation Details, we elaborate on the implementation details on the GPT-4 prompts for name generation, human verification process for RENOVATE names on evaluation sets, open metrics, and the renaming and pre-trained models.
@@
  • In Appendix D, More Name Examples, we present more examples of context, candidate, and renovated names. We also show more examples of name selection with the IoU scores.
@@
  • In Appendix E, More Qualitative Analysis, we provide more qualitative analysis demonstrating that RENOVATE can refine original names, correct wrong annotations, and uncover segments with shared semantic concepts across datasets.
  • In Appendix F, License for Existing Assets, we list the names of the license for each assets used in this paper.
Appendix A More Literature Review
@@
In this section, we provide a brief overview of name problems in open-vocabulary classification. While OpenSeg [5] is the only prior work for the name problems in open-vocabulary segmentation, there are more research works on improving name quality for open-vocabulary classification. For example, [48] improve class names by including their parents and children from the WordNet hierarchy [49].  [30] propose to replace the fixed vocabulary list with nouns extracted from best-matching captions.  [50] improve the descriptiveness of the vocabulary by adding class descriptions. However, since the VLMs used by these models are all trained on image-text datasets, it is non-trivial to extend them to dense prediction tasks like segmentation. For the renaming strategy, our context name design and LLM-based generation approach ensure the diversity of our names with hierarchical concepts as well as descriptive contexts, without the use of an external word hierarchy as [49] or class descriptions as [50]. Our renaming model further improves the name quality based on the visual segments and is more specifically designed for the segmentation tasks.
Appendix B More Experiment Results
(a) User interface for human preference study.
(b) Human preference results.
Figure B.1: Human preference study. A survey of 20 researchers is conducted to compare preferences between the original names versus RENOVATE names on the validation sets. RENOVATE names are favored in 82% cases.
B.1 Human preference study
As demonstrated in Fig. B.1, we study the name quality of our renovated names with a human preference study on ADE20K and Cityscapes datasets, which consists of 30249 and 14670 segments in the validation set, respectively. Specifically, we first randomly select 10% of all the segments in the validation set of each dataset. The average time for human experts to finish the preference study is about 2 hours. Then for each segment, we ask 20 researchers unfamiliar with the project about their preference between its original name and our model-selected name. Our results demonstrate that our renovated names are preferred in 82% cases, showing a clear advantage over the original names. We provide more details of our human study in the supplements.
B.2 Ablation analysis on the renaming pipeline
To conduct ablation study on the renaming pipeline, we leverage pre-trained open-vocabulary model models as evaluators of names, automating the name quality evaluation process. Instead of involving humans, we prompt pre-trained open-vocabulary segmentation models with different names. Names are then considered to be better if they help the model achieve better open-vocabulary segmentation. Specifically, we use pre-trained FC-CLIP model for the following experiments.
In Table B.2, we first analyze the impact of context names for candidate generation. Our results indicate significant performance improvements when context information is provided, as opposed to the “None” approach, which relies only on original names. This highlights the effectiveness of leveraging contextual insights for generating superior candidate names. In comparison, CaSED outperforms both BLIP2 and RAM, achieving a greater boost in performance through its retrieval-based approach and the utilization of expansive external image-caption datasets, underlining the benefits of contextual enrichment in name generation.
ADE20K
Context name source PQ AP mIoU
None 25.50 16.53 33.80
Caption nouns (BLIP2 [51]) 25.89 17.36 34.74
Image tags (RAM [52]) 26.95 17.55 35.54
Caption nouns (CaSED [30]) 27.90 17.88 37.05
Table B.1: Ablation on the context name sources.
ADE20K
Attn bias Query Classifier PQ AP mIoU
gt mask learnable text 21.09 14.91 26.96
gt mask text learnable 25.65 16.33 32.67
rand gt mask text learnable 26.21 17.50 35.62
rand gt mask text w/ neg learnable 27.90 17.88 37.05
Table B.2: Ablation on renaming model architecture.
In Table B.2, we analyze the design choices of our renaming model. We start by establishing that our approach using text queries is superior to a standard text classifier, which only interacts with the transformer decoder’s output query embeddings and treats the mask prediction branch in a class-agnostic way. In comparison, our text queries interact intimately with spatial details from image features and exploit the mask prediction branch, leading to more precise name-segment matching. We also observe performance benefits from randomly substituting ground-truth masks with predicted ones in attention biases (“rand gt mask”), which suggests training with increased reliance on text queries refines the matching quality. Furthermore, augmenting our model with negative name appending (“text w/ neg”) enhances its discriminative capacity, further boosting performance and solidifying the renaming model’s overall effectiveness.
ADE20K [14] Cityscapes [15]
Model Benchmark Names Metric PQ AP mIoU PQ AP mIoU
ODISE Original Standard 21.88 13.94 29.16 39.72 27.73 49.60
Open 26.31 14.49 34.39 44.09 27.91 58.20
OpenSeg Standard 21.63 13.95 28.88 43.26 28.45 54.53
Open 28.02 15.32 38.86 46.66 28.88 58.61
RENOVATE Standard 23.69 14.38 31.64 43.61 28.38 57.42
Open 42.32 50.11 44.65 53.10 34.52 65.05
MasQCLIP Original Standard 23.46 12.80 30.32 33.78 18.11 45.35
Open 26.53 13.01 36.33 37.94 18.35 55.75
OpenSeg Standard 23.70 12.84 31.17 35.05 17.79 46.46
Open 28.04 13.89 39.40 38.83 18.62 52.11
RENOVATE Standard 25.00 12.93 32.30 35.63 18.19 50.07
Open 47.59 67.01 42.75 49.30 37.25 61.44
FC-CLIP Original Standard 24.30 16.79 32.14 39.42 22.76 53.08
Open 27.04 16.52 36.59 42.57 26.54 62.69
OpenSeg Standard 25.35 17.30 33.06 43.68 26.93 56.15
Open 29.03 17.52 41.03 45.92 27.09 60.21
RENOVATE Standard 27.90 17.88 37.05 45.90 29.79 62.55
Open 58.25 66.99 48.31 58.63 38.91 70.60
Table B.3: Open-vocabulary segmentation on using both standard and open metrics across different pre-trained models using different name sets.
B.3 Improving evaluation benchmarks with renovated names
In Table B.3, we show the full evaluation results of the experiments described in Section 5.3. With results on ADE20K and Cityscapes with PQ, AP, and mIoU evaluation, we make similar observations as Section 5.3 that the performance gap between standard and open metrics is significantly larger when evaluating using RENOVATE names, indicating that many misclassification cases are indeed “benign”. We also provide more examples of fine-grained misclassification analysis in Fig. B.2.
Figure B.2: More examples of fine-grained misclassification analysis.
Appendix C More Implementation Details
C.1 GPT-4 prompts for name generation
@@
For the system message, we start the message with “You are a helpful assistant” following the examples in OpenAI Chat Completions API [53]. We then describe our problem setup, input format, and instructions with examples on name generation. For the user message, {original class name} and {context names} require inputs from each category. For our ablation experiments without context names (“None” in Table B.2 in the main paper), we remove the descriptions about context names in the system message and do not require context names in the user message.
C.2 More details on the human verification process
To ensure the high quality of names on the evaluation sets of ADE20K and Cityscapes, we verify the names on all the segments. As shown in Fig. 1(a), for each segment, we provide the model-selected name along with the other top 3 names ranked by the model, we also provide the rest of the candidate names in the “others” option. Therefore, the human verifier can quickly continue to the next segment if the selected name itself or another name in top 3 names matches the most with the segment. If none of the top 3 names fits well, humans need to choose from the other candidate names in “others”. In the latter case, it would be similar to the scenario where RENOVATE suggestions are not provided and humans have to choose from the whole un-ranked candidate name list. In Fig. 1(b), we compare the name verification speed with and without RENOVATE suggestions on 200 segments. As we can see, RENOVATE suggestions significantly speed up our verification process. We note that the human verifiers are provided only with candidate names from the same original class, which is typically the original ground-truth class. In the rare cases where our renaming model prefers names from classes other than the original groundtruth class, we first verify the class choices, then provide the human verifiers with the name suggestions from the verified class.
We further note that for all the segments verified, 68% of them are matched to the default, i.e., top 1, model-selected name, and 92% of them are matched to the top 3 RENOVATE suggestions, reflecting the high quality of RENOVATE names. In Fig. C.2, we show typical cases when humans and models disagree, including segments that are too small to recognize and segments whose names cannot be inferred correctly due to insufficient visual cues. We note that such segments only consist a minority of all the segments, with inherent difficulty to name precisely.